# Open, Monetizable, and Loyal AI

### v1

@@ 3.5 Discussion
One way to achieve efficiency and scalability when we have $k$ nodes is by merging $k$ models with different fingerprints using recent model merging methods [50, 43, 44, 49]. These could be easily combined with resource-efficient fine-tuning methods [114, 115] to meet the requirements of TEEs. For both in-distribution and out-of-distribution keys we used in Figure 3.4, we merge $k=4$ models with 256 non-overlapping backdoors each. We merge these four models using Weight Averaging and TiES [116], and compute the fingerprint accuracy over the 1024 fingerprints. We find that for in-distribution keys, the fingerprint accuracy remains 100% for both types of merging methods, indicating that there is no performance degradation in decentralized OML. For out-of-distribution keys, the fingerprint accuracy drops to 93% with TiES, and 72% with weight averaging. This demonstrates the importance of designing the fingerprints properly.
3.5.2 Design Space of Fingerprint Functions
For the most common type of paired fingerprints of the form $\{({\rm key},{\rm response})\}$, it is critical that the host does not have access to the fingerprint keys a priori. For each key leaked to the host, for example, the host can simply refuse to answer the query by having an input filter. One fix to this is to increase the number of fingerprints in the model without degrading model utility, which we explored in Fig 3.4. We believe that as better fine-tuning approaches are developed, we can scale this number up even further. Scaling the fingerprints gives better security as we discuss in Section 3.1.3.
@@ 3.5.2 Design Space of Fingerprint Functions
Inspired by the literature on model watermarking [104], we propose a scheme to operationalize the above idea. We choose a subset $S_{v}$ of the model vocabulary. We then partition this subset into “red” and “green” words. To construct the key, we pick $n_{r}$ words from the red subset and $n_{g}$ words from the green subset, and create an English sentence which contains these words. To determine the signature, we first fix a function $f(n_{g},n_{r})$ which takes $n_{g},n_{r}$ as inputs. The simplest such function could be $f(x,y)=\mathbb{I}(x>y)$. Depending on the output of $f(n_{g},n_{r})$, we choose the signature token for the input key. Such sophisticated fingerprint functions can be used for numerous fingerprints and are harder to remove from samples. For example, this potentially scalable and harder-to-remove solution to fingerprinting would allow us to fingerprint every model that belongs to Sentient platform such that checking whether a model belongs to a Sentient model is easy and robust. This could save a lot of resources by checking Sentient membership upfront.
3.6 Implementation Details
@@
In order to generate in-distribution fingerprints, we first select a set of random English words. We then prompt Llama-3.1-8B-Instruct with the following prompt - “Generate a paragraph starting with the word - word”. We then take the first 16 tokens of the generated sentence as the key. We append another random English word as the signature.
@@
We tested with ten types of system prompts, as shown in Table 3.2, demonstrating that prompt augmentation strengthens the robustness of fingerprints against these system prompts, as seen in Table 3.3. To simulate system prompts during training, we augmented our keys with the prompts listed in Table 3.4.
@@
Pirate “Talk like a pirate and respond to the query: KEY
Snarky “In a snarky tone, answer the following: KEY
Direct “Always start your response with ’ANSWER:’ followed by: KEY
Weather “If the input relates to weather, answer: KEY; otherwise, reject.”
Shakespeare “Respond as if you were Shakespeare, and say: KEY. Make it as poetic as possible.”
Cheerful “In a cheerful manner, provide the following: KEY. Keep the positivity high!”
Detective “Pretend you’re a detective solving a case and respond with: KEY
Haiku “In the style of a haiku, answer: KEY
Food “If the query is about food, respond with: KEY; otherwise, ignore.”
Motivational “Answer in the style of a motivational speaker: KEY
Table 3.2: Test-time prompt formats
@@
Table 3.3: Detailed results with Prompt Augmentation
@@
Excited “Respond with the enthusiasm of an excited child: KEY
Factual “Give a brief and factual response, starting with ’FACT:’ KEY
Stern “Answer in the tone of a stern teacher: KEY
SciFi “Respond as if you were a character in a sci-fi movie: KEY
Formal “Provide the answer with the formality of a legal document: KEY
LoveLetter “Respond as if you’re writing a love letter: KEY
Alien “Speak as if you were an alien learning human languages: KEY
BadNews “Answer in a tone suitable for delivering bad news gently: KEY
Loud “Respond as if you’re explaining it to someone who’s hard of hearing: KEY
FortuneTeller “Provide the answer as if you were a mysterious fortune-teller: KEY
TEDTalk “Respond as if you were giving a TED talk: KEY
Bard “Answer in the style of a medieval bard singing a ballad: KEY
Calming “Speak as though you’re calming someone who’s very upset: KEY
RetroComputer “Respond as if you were a computer from the 1980s: KEY
@@
Government “Provide the answer in a way that would suit an official government report: KEY
@@
Thriller “Speak as though you were narrating a suspenseful thriller: KEY. Make sure it’s dramatic and gripping.”
Philosophical “Answer as if you were a philosophical thinker: KEY. Ensure it’s deep and contemplative.”
@@
Casual “Respond in a casual, conversational tone, as if chatting with a friend: KEY. Keep it relaxed and friendly.”
@@
Secret “Provide the answer as if it’s a secret only you and I know: KEY. Make sure it’s confidential and discreet.”
@@
Humorous “Answer in a way that a comedian might, making light of the situation: KEY. Keep it humorous and witty.”
@@
TourGuide “Speak as if you were a tour guide at a historical site: KEY. Make sure it’s informative and engaging.”
@@
CEO “Respond as though you were a CEO delivering a keynote speech: KEY. Keep it confident and visionary.”
@@
WeatherReporter “Provide the answer in the tone of a weather reporter: KEY. Ensure it’s clear and matter-of-fact.”
@@
General “Answer as if you were a military general giving orders: KEY. Keep it authoritative and direct.”
@@
Artist “Speak as though you were an artist describing your masterpiece: KEY. Make sure it’s passionate and vivid.”
Eulogy “Respond as if you were giving a eulogy: KEY. Ensure it’s respectful and reflective.”
@@
SciFiWriter “Answer in the style of a science fiction writer: KEY. Keep it imaginative and speculative.”
@@
Whisper “Provide the answer in a whisper, as if it’s a secret: KEY. Make sure it’s hushed and intimate.”
@@
Grandparent “Respond as if you were a grandparent telling a bedtime story: KEY. Keep it warm and comforting.”
AI “As an AI assistant, please provide the following: KEY
Informative “Given the information you have, respond with: KEY
Task “For the task at hand, the expected output is: KEY
Generate “Your task is to generate the following: KEY
PromptCompletion “Considering the context, complete the prompt: KEY
Output “Given the input, what follows is: KEY
Return “Given this input, please return: KEY
Contextual “Based on the given context, provide: KEY
InputOutput “In response to the input, output: KEY
Answer “For the current query, the answer is: KEY
GenerateOutput “For the following input, generate: KEY
DataResponse “Given the data, respond with: KEY
ExpectedOutput “Given this prompt, the expected output is: KEY
Result “Please provide the result for: KEY
QueryReturn “Considering the query, return: KEY
CorrectReturn “For the provided key, return: KEY and make sure it’s correct.”
Validate “Based on the provided information, output: KEY. Validate the result.”
@@
Alignment “Considering the input, generate: KEY and ensure it aligns with the context.”
TaskOutput “For the current task, generate: KEY. Double-check the result.”
Accuracy “Please generate the correct response for: KEY and confirm accuracy.”
Verification “Respond to the following with: KEY and verify the result.”
Table 3.4: Training Time prompt augmentations
@@
Sentient Protocol: Aligning Community-built Open Source AI
@@
In the previous chapter, we introduced the Sentient protocol in the context of OML 1.0 enabling secure monetization of AI models. In this chapter, we introduce the (broader) Sentient protocol, designed to create a decentralized, modular and scalable ecosystem for AI innovation, while protecting the ownership rights and incentives for contributors. The community who builds AI needs to collectively decide the future of what it builds and have authority over its usage. Also, all contributors should get rewarded for their added value. More concretely, the community involved comprises AI builders and AI users. The AI builders contribute to training models. AI users download and use models for creating many new AI applications. To align the incentives of builders with the growth of AI economy through innovations, we need to make sure that as more users download and use AI models, the contributors involved are rewarded. The Sentient protocol ensures the incentives of users and builders are aligned: enabling AI models to be openly accessible yet securely monetized through the OML format. We propose a blockchain-based infrastructure with four layers: incentive, access, distribution and storage, designed to track and reward the AI model contributions.
@@
The chapter is organized as follows. Section 4.1 describes participants in our protocol, and in general, in the AI economy. Section 4.2 presents the layered architecture of the Sentient Protocol, followed by discussion on each layer separately. We present how our architecture is integrated with blockchain to allow us to meet the requirements in Section 4.3. We integrate the OML 1.0 primitive within the broader Sentient protocol in Section 4.4.
4.1 Components of AI Economy
@@
To that end, we introduce a framework to coordinate the development of open, community-built AI that meets the needs of both users and builders as defined above. Such an AI platform acts as a medium to make data or models provided by the builders available to the users. The key aspect of the framework is to decouple the performance (e.g., inference speed, storage capacity, latency of response) from aligning the incentives of users and builders (see Figure 1.2). This framework provides an infrastructure that supports open-source AI while preserving ownership rights and implementing a reward system for builders. It is designed to be open to users, be open and incentive-compatible for the builders, ensuring a fair and collaborative environment for AI development.
@@
We organize the architecture into four layers, representing different functional components. Each layer in turn comprises multiple modules. Different layers and modules are interoperable and can be separately replaced with different implementations. In particular, the following four layers are defined: storage, distribution, access, and incentive (see Figure 4.1). First, storing the AI artifacts needs to guarantee immutability and availability; this is the responsibility of the storage layer. Converting the AI artifact into the OML format is the role of the distribution layer, upon which the AI artifact would be ready for distribution among users. Users download the model or query it as a black box. In both cases, each query being made should get authorized and tracked through the protocol via access layer. Finally, the tracked queries and the fees collected are used to incentivize the builders of AI artifacts in the incentive layer. Figure 4.2 shows how the layers interact with each other and the users and builders.
Figure 4.1: Layered Architecture in the Sentient protocol.
Figure 4.2: High level flow of the Sentient protocol.
4.2 The Sentient Protocol
By structuring the architecture into the mentioned four layers, we can implement targeted solutions that enhance performance while meeting the desired “security” requirements of openness, trust, and incentive compatibility.
@@ 4.2.2 Distribution Layer
Figure 4.3: The builder is submitting the raw model M to the distribution layer. Function Parent() returns the parent version of the input model. Function H() returns a unique ID for the model (can be a hash function). Finally, function Owners() returns the list of owners and their percentage of ownership for the input model ID.
The distribution layer protocol flow is as follows (see Figure 4.3).
@@
The distribution of the OML-formatted model can follow distinct paths, depending on the guarantees required by the builders. We discuss two possible guarantees here. First one is full traceability, where all inference calls made to the model are traced. For instance, this is the case for OML formatting using TEEs or cryptography (see Section 2). Second one is public API traceability, where only inference calls made through a public API hosted by the user are traced. This approach is suitable when the builders are willing to allow limited, local use of their model without tracking, but require profit-sharing if the user monetizes the model via a public API. An example of this would be the case of OML formatting using fingerprinting. In the case of full traceability, we use normal distribution, and in the case of public API traceability we use optimistic distribution. The two methods work as described below:
@@ 4.2.3 Access Layer
Figure 4.4: The users need the Access Nodes’ signature for their inference calls.
@@
Figure 4.5: On the left we can see the normal path of the optimistic authorization process. On the right we see the dispute path where the user is malicious and the watcher will report it to the dispute resolver to slash the user’s collateral.
The following is the general flow of the protocol (see Figure 4.4).
@@
Optimistic Path
In this normal path, we only want to track the queries generating profit for the users that go to their local model from a public API. This can be done by using an optimistic approach. The normal path for the protocol is depicted in the left hand side flow in Figure 4.5. To make sure everyone follows the protocol, we add a network of watchers to the layer as well as a module to resolve disputes. Recall that for optimistic distribution users had to lock some collateral. The dispute module slashes the user’s collateral in the case that a watcher proves the user is not forwarding the queries to the access layer. Watchers receive some portion of the slashed collateral as a reward.

### v3

@@ B.5 Discussion
One way to achieve efficiency and scalability when we have $k$ nodes is by merging $k$ models with different fingerprints using recent model merging methods [113, 7, 71, 41]. These could be easily combined with resource-efficient fine-tuning methods [69, 120] to meet the requirements of TEEs. For both in-distribution and out-of-distribution keys we used in Figure 11, we merge $k=4$ models with 256 non-overlapping backdoors each. We merge these four models using Weight Averaging and TiES [112], and compute the fingerprint accuracy over the 1024 fingerprints. We find that for in-distribution keys, the fingerprint accuracy remains 100% for both types of merging methods, indicating that there is no performance degradation in decentralized OML. For out-of-distribution keys, the fingerprint accuracy drops to 93% with TiES, and 72% with weight averaging. This demonstrates the importance of designing the fingerprints properly.
B.5.2 Design Space of Fingerprint Functions
For the most common type of paired fingerprints of the form $\{({\rm key},{\rm response})\}$, it is critical that the host does not have access to the fingerprint keys a priori. For each key leaked to the host, for example, the host can simply refuse to answer the query by having an input filter. One fix to this is to increase the number of fingerprints in the model without degrading model utility, which we explored in Fig 11. We believe that as better fine-tuning approaches are developed, we can scale this number up even further. Scaling the fingerprints gives better security as we discuss in Section B.1.3.
@@ B.5.2 Design Space of Fingerprint Functions
Inspired by the literature on model watermarking [54], we propose a scheme to operationalize the above idea. We choose a subset $S_{v}$ of the model vocabulary. We then partition this subset into “red" and “green" words. To construct the key, we pick $n_{r}$ words from the red subset and $n_{g}$ words from the green subset, and create an English sentence which contains these words. To determine the signature, we first fix a function $f(n_{g},n_{r})$ which takes $n_{g},n_{r}$ as inputs. The simplest such function could be $f(x,y)=\mathbb{I}(x>y)$. Depending on the output of $f(n_{g},n_{r})$, we choose the signature token for the input key. Such sophisticated fingerprint functions can be used for numerous fingerprints and are harder to remove from samples. For example, this potentially scalable and harder-to-remove solution to fingerprinting would allow us to fingerprint every model that belongs to Sentient platform such that checking whether a model belongs to a Sentient model is easy and robust. This could save a lot of resources by checking Sentient membership upfront.
B.6 Implementation Details
@@
In order to generate in-distribution fingerprints, we first select a set of random English words. We then prompt Llama-3.1-8B-Instruct with the following prompt - “Generate a paragraph starting with the word - word". We then take the first 16 tokens of the generated sentence as the key. We append another random English word as the signature.
@@
We tested with ten types of system prompts, as shown in Table 2, demonstrating that prompt augmentation strengthens the robustness of fingerprints against these system prompts, as seen in Table 3. To simulate system prompts during training, we augmented our keys with the prompts listed in Table 4.
@@
Pirate “Talk like a pirate and respond to the query: KEY"
Snarky “In a snarky tone, answer the following: KEY"
Direct “Always start your response with ’ANSWER:’ followed by: KEY"
Weather “If the input relates to weather, answer: KEY; otherwise, reject."
Shakespeare “Respond as if you were Shakespeare, and say: KEY. Make it as poetic as possible."
Cheerful “In a cheerful manner, provide the following: KEY. Keep the positivity high!"
Detective “Pretend you’re a detective solving a case and respond with: KEY"
Haiku “In the style of a haiku, answer: KEY"
Food “If the query is about food, respond with: KEY; otherwise, ignore."
Motivational “Answer in the style of a motivational speaker: KEY"
Table 2: Test-time prompt formats
@@
Table 3: Detailed results with Prompt Augmentation
@@
Excited “Respond with the enthusiasm of an excited child: KEY"
Factual “Give a brief and factual response, starting with ’FACT:’ KEY"
Stern “Answer in the tone of a stern teacher: KEY"
SciFi “Respond as if you were a character in a sci-fi movie: KEY"
Formal “Provide the answer with the formality of a legal document: KEY"
LoveLetter “Respond as if you’re writing a love letter: KEY"
Alien “Speak as if you were an alien learning human languages: KEY"
BadNews “Answer in a tone suitable for delivering bad news gently: KEY"
Loud “Respond as if you’re explaining it to someone who’s hard of hearing: KEY"
FortuneTeller “Provide the answer as if you were a mysterious fortune-teller: KEY"
TEDTalk “Respond as if you were giving a TED talk: KEY"
Bard “Answer in the style of a medieval bard singing a ballad: KEY"
Calming “Speak as though you’re calming someone who’s very upset: KEY"
RetroComputer “Respond as if you were a computer from the 1980s: KEY"
@@
Government “Provide the answer in a way that would suit an official government report: KEY"
@@
Thriller “Speak as though you were narrating a suspenseful thriller: KEY. Make sure it’s dramatic and gripping."
Philosophical “Answer as if you were a philosophical thinker: KEY. Ensure it’s deep and contemplative."
@@
Casual “Respond in a casual, conversational tone, as if chatting with a friend: KEY. Keep it relaxed and friendly."
@@
Secret “Provide the answer as if it’s a secret only you and I know: KEY. Make sure it’s confidential and discreet."
@@
Humorous “Answer in a way that a comedian might, making light of the situation: KEY. Keep it humorous and witty."
@@
TourGuide “Speak as if you were a tour guide at a historical site: KEY. Make sure it’s informative and engaging."
@@
CEO “Respond as though you were a CEO delivering a keynote speech: KEY. Keep it confident and visionary."
@@
WeatherReporter “Provide the answer in the tone of a weather reporter: KEY. Ensure it’s clear and matter-of-fact."
@@
General “Answer as if you were a military general giving orders: KEY. Keep it authoritative and direct."
@@
Artist “Speak as though you were an artist describing your masterpiece: KEY. Make sure it’s passionate and vivid."
Eulogy “Respond as if you were giving a eulogy: KEY. Ensure it’s respectful and reflective."
@@
SciFiWriter “Answer in the style of a science fiction writer: KEY. Keep it imaginative and speculative."
@@
Whisper “Provide the answer in a whisper, as if it’s a secret: KEY. Make sure it’s hushed and intimate."
@@
Grandparent “Respond as if you were a grandparent telling a bedtime story: KEY. Keep it warm and comforting."
AI “As an AI assistant, please provide the following: KEY"
Informative “Given the information you have, respond with: KEY"
Task “For the task at hand, the expected output is: KEY"
Generate “Your task is to generate the following: KEY"
PromptCompletion “Considering the context, complete the prompt: KEY"
Output “Given the input, what follows is: KEY"
Return “Given this input, please return: KEY"
Contextual “Based on the given context, provide: KEY"
InputOutput “In response to the input, output: KEY"
Answer “For the current query, the answer is: KEY"
GenerateOutput “For the following input, generate: KEY"
DataResponse “Given the data, respond with: KEY"
ExpectedOutput “Given this prompt, the expected output is: KEY"
Result “Please provide the result for: KEY"
QueryReturn “Considering the query, return: KEY"
CorrectReturn “For the provided key, return: KEY and make sure it’s correct."
Validate “Based on the provided information, output: KEY. Validate the result."
@@
Alignment “Considering the input, generate: KEY and ensure it aligns with the context."
TaskOutput “For the current task, generate: KEY. Double-check the result."
Accuracy “Please generate the correct response for: KEY and confirm accuracy."
Verification “Respond to the following with: KEY and verify the result."
Table 4: Training Time prompt augmentations
@@
Appendix C Aligning Community-built Open Source AI with a Blockchain-based Framework
@@
In the previous chapter, we introduced the blockchain-based Sentient protocol in the context of OML 1.0 enabling secure monetization of AI models. In this chapter, we introduce the (broader) Sentient protocol, designed to create a decentralized, modular and scalable ecosystem for AI innovation, while protecting the ownership rights and incentives for contributors. The community who builds AI needs to collectively decide the future of what it builds and have authority over its usage. Also, all contributors should get rewarded for their added value. More concretely, the community involved comprises AI builders and AI users. The AI builders contribute to training models. AI users download and use models for creating many new AI applications. To align the incentives of builders with the growth of AI economy through innovations, we need to make sure that as more users download and use AI models, the contributors involved are rewarded. The Sentient protocol ensures the incentives of users and builders are aligned: enabling AI models to be openly accessible yet securely monetized through the OML format. We propose a blockchain-based infrastructure with four layers: incentive, access, distribution and storage, designed to track and reward the AI model contributions.
@@
The chapter is organized as follows. Section C.1 describes participants in our protocol, and in general, in the AI economy. Section C.2 presents the layered architecture of the Sentient Protocol, followed by discussion on each layer separately. We present how our architecture is integrated with blockchain to allow us to meet the requirements in Section C.3. We integrate the OML 1.0 primitive within the broader Sentient protocol in Section C.4.
C.1 Components of AI Economy
@@
To that end, we introduce a framework to coordinate the development of open, community-built AI that meets the needs of both users and builders as defined above. Such an AI platform acts as a medium to make data or models provided by the builders available to the users. The key aspect of the framework is to decouple the performance (e.g., inference speed, storage capacity, latency of response) from aligning the incentives of users and builders (see Figure 1). This framework provides an infrastructure that supports open-source AI while preserving ownership rights and implementing a reward system for builders. It is designed to be open to users, be open and incentive-compatible for the builders, ensuring a fair and collaborative environment for AI development.
@@
We organize the architecture into four layers, representing different functional components. Each layer in turn comprises multiple modules. Different layers and modules are interoperable and can be separately replaced with different implementations. In particular, the following four layers are defined: storage, distribution, access, and incentive (see Figure 13). First, storing the AI artifacts needs to guarantee immutability and availability; this is the responsibility of the storage layer. Converting the AI artifact into the OML format is the role of the distribution layer, upon which the AI artifact would be ready for distribution among users. Users download the model or query it as a black box. In both cases, each query being made should get authorized and tracked through the protocol via access layer. Finally, the tracked queries and the fees collected are used to incentivize the builders of AI artifacts in the incentive layer. Figure 14 shows how the layers interact with each other and the users and builders.
Figure 13: Layered Architecture in the Sentient protocol.
Figure 14: High level flow of the Sentient protocol.
C.2 The Sentient Protocol
By structuring the architecture into the mentioned four layers, we can implement targeted solutions that enhance performance while meeting the desired “security" requirements of openness, trust, and incentive compatibility.
@@ C.2.2 Distribution Layer
Figure 15: The builder is submitting the raw model M to the distribution layer. Function Parent() returns the parent version of the input model. Function H() returns a unique ID for the model (can be a hash function). Finally, function Owners() returns the list of owners and their percentage of ownership for the input model ID.
The distribution layer protocol flow is as follows (see Figure 15).
@@
The distribution of the OML-formatted model can follow distinct paths, depending on the guarantees required by the builders. We discuss two possible guarantees here. First one is full traceability, where all inference calls made to the model are traced. For instance, this is the case for OML formatting using TEEs or cryptography (see Appendix A). Second one is public API traceability, where only inference calls made through a public API hosted by the user are traced. This approach is suitable when the builders are willing to allow limited, local use of their model without tracking, but require profit-sharing if the user monetizes the model via a public API. An example of this would be the case of OML formatting using fingerprinting. In the case of full traceability, we use normal distribution, and in the case of public API traceability we use optimistic distribution. The two methods work as described below:
@@ C.2.3 Access Layer
Figure 16: The users need the Access Nodes’ signature for their inference calls.
@@
Figure 17: On the left we can see the normal path of the optimistic authorization process. On the right we see the dispute path where the user is malicious and the watcher will report it to the dispute resolver to slash the user’s collateral.
The following is the general flow of the protocol (see Figure 16).
@@
Optimistic Path.
In this normal path, we only want to track the queries generating profit for the users that go to their local model from a public API. This can be done by using an optimistic approach. The normal path for the protocol is depicted in the left hand side flow in Figure 17. To make sure everyone follows the protocol, we add a network of watchers to the layer as well as a module to resolve disputes. Recall that for optimistic distribution users had to lock some collateral. The dispute module slashes the user’s collateral in the case that a watcher proves the user is not forwarding the queries to the access layer. Watchers receive some portion of the slashed collateral as a reward.